A COMPLETE STRUCTURE FOR CIS,TRANS-1,4-DIFLUOROBUTADIENE FROM MICROWAVE SPEC-TROSCOPY

\underline{NORMAN C. CRAIG}, CATHERINE M. OERTEL, and DAVID C. OERTEL, Department of Chemistry, Oberlin College, Oberlin, OH 44074; MICHAEL J. TUBERGEN and RICHARD J. LAVRICH, Department of Chemistry, Kent State University, Kent, OH 44242.; ANNE M. CHAKA, The Lubrizol Corporation, Wickliffe, OH 44092-2298.

Microwave spectra have been observed for cis, trans-1,4-difluorobutadiene, for the four ^{13}C isotopomers in natural abundance, and for five deuterium isotopomers in a mixture prepared by partial exchange in NaOD/D_{2}O. Watson-type Hamiltonians have been fitted to the a-type and b-type transitions observed for each species. For the normal species, which is a near prolate top, A = 12988.3327(11), B = 1467.8791(3), and C = 1318.5845(3) MHz. From Stark effect measurements, the dipole moment components are μ_{a} = 0.660(4) D and μ_{b} = 2.213(5) D. A Kraitchman substitution analysis succeeds for the carbon atom backbone and for the hydrogen atoms on the end carbon atoms but does not for the interior hydrogen atoms. Most notably, I_{b}’ for the 2-d_{1} species is smaller than I_{b} for the normal species. To obtain a complete structure, results of DFT calculations with the adiabatic connection method were used for the bond lengths and bond angles of the interior C-H bonds. Otherwise the agreement of experiment and theory is good.